Ballery1:

Find the solution set of the following equations: X-2y+3z = 5

jhonyy9:

,,equations" - where are ?

mhchen:

It looks like there's only 1 equation, and 3 variables. x = 2y - 3z + 5 y = y z = z \[\left[\begin{matrix}x \\ y \\ z\end{matrix}\right] = \left[\begin{matrix}2r -3s + 5\\ r \\ s\end{matrix}\right] = r\left[\begin{matrix}2\\ 1 \\0 \end{matrix}\right] + s \left[\begin{matrix}-3\\ 0 \\ 1\end{matrix}\right] + \left[\begin{matrix}5 \\ 0 \\ 0\end{matrix}\right]\] r and s can be anything. Are you sure your question is right though? xd

imqwerty:

This equation represents a plane, you can represent the solutions with the parametric form

Ballery1:

Yup, there’s only one equation with three variables.

Ballery1:

Let me post the actual picture of the question for you guys lol

mhchen:

if it's just 1 equation, then you have your answer right up there.

Ballery1:

Ballery1:

The answer is written in some other way... let me post ..one sec

Ballery1:

Ballery1:

Wait, let me show you a method that I’ve been using to solve these questions... one second plz

mhchen:

So their k1 and k2 is just like my r and s. But I think they did z = (5/3) + (2/3)y - (1/3)x

Ballery1:

So this is the template I’ve been using to solve these kinda questions

mhchen:

Yeah that works

Ballery1:

But please tell me why did they divide both sides by 3 instead of 2 or 1??

Ballery1:

Do you want me to draw for you sir? :)

mhchen:

because they were solving for z

Ballery1:

Why z in particular?

mhchen:

it doesn't have to be z, it can be x or y actually.

Ballery1:

So does that mean if you divide and isolate for let’s say y variable, we should come to the same answer in the end?

Ballery1:



mhchen:

So my solution was \[\left[\begin{matrix}x \\ y \\ z\end{matrix}\right] = \left[\begin{matrix}2r -3s + 5\\ r \\ s\end{matrix}\right] \] and their solution was \[\left[\begin{matrix}x \\ y \\ z\end{matrix}\right] = \left[\begin{matrix}r\\ s \\ \frac{5}{3}-\frac{1}{3}r+\frac{2}{3}s\end{matrix}\right] \] hmmm

mhchen:

I never explored this, but theoretically...the should have the same "Solution space" There's infinite number of solutions, but not every number is a solution, so we call it a solution space.

Ballery1:

Is there a video on this topic ?

Ballery1:

My textbook literally didn’t do any sample questions for this particular topic

mhchen:

3blue1brown might have a good video on this

Ballery1:

Thanks bro, really appreciated:)